# Deep Context-Conditioned Anomaly Detection for Tabular Data

### v1

@@ Abstract.
Anomaly detection is critical in domains such as cybersecurity and finance, especially when working with large-scale tabular data. Yet, unsupervised anomaly detection—where no labeled anomalies are available—remains a significant challenge. Although various deep learning methods have been proposed to model a dataset’s joint distribution, real-world tabular data often contain heterogeneous contexts (e.g., different users), making globally rare events normal under certain contexts. Consequently, relying on a single global distribution can overlook these contextual nuances, degrading detection performance. In this paper, we present a context-conditional anomaly detection framework tailored for tabular datasets. Our approach automatically identifies context features and models the conditional data distribution using a simple deep autoencoder. Extensive experiments on multiple tabular benchmark datasets demonstrate that our method outperforms state-of-the-art approaches, underscoring the importance of context in accurately distinguishing anomalous from normal instances.
Unsupervised Anomaly Detection, Deep Autoencoder
$c$s: Computing methodologies Anomaly detection
1. Introduction
@@
Unsupervised anomaly detection (AD) is a critical field focused on identifying patterns that deviate from expected behavior, with applications spanning a wide range of tasks like security, fraud detection, and system reliability (Chandola et al., 2009; Aggarwal, 2017). The core challenge is not merely detecting new or unseen objects, but discerning whether such observations align with the learned notion of “normal,” or truly represent anomalies.
@@
Recent advances in deep learning have driven a wave of neural network-based methods for unsupervised anomaly detection (Landauer et al., 2023; Pang et al., 2021b). Despite methodological differences, most deep approaches model the joint distribution of all features, implicitly assuming that data patterns are homogeneous across conditions. In practice, patterns often vary significantly with context. For example, in cybersecurity, a spike in data transfer may be benign for a user who routinely performs backups but suspicious for a database administrator. Ignoring such context can degrade detection accuracy.
@@
To address this challenge, we argue probabilistically and demonstrate empirically the effectiveness of modeling conditional distributions over joint distributions for anomaly detection. We advocate for the context-conditional anomaly detection framework, which models conditional distributions to capture intra-context regularities while isolating inter-context variability, thereby enhancing detection performance without requiring separate models for each condition.
@@
Selecting an appropriate context feature is a critical aspect of contextual learning. It is often not obvious what feature one should condition on when modeling the data. Indeed, the choice of conditioning variable can significantly influence model performance. To address this, we develop a bi-level optimization methodology to automatically identify the context features based on negative log-likelihood. Requiring only a small set of anomaly-free validation samples, our approach is simple, model-agnostic, and demonstrates consistent benefits across a wide range of datasets.
@@
To empirically validate our claim, we leverage recent advancements in deep learning and advocate for the conditional Wasserstein autoencoder (CWAE), a conditional variant of the Wasserstein autoencoder (WAE) (Tolstikhin et al., 2019), to model conditional distributions for anomaly detection. We conduct experiments using several tabular AD benchmark datasets containing inherent contextual information (e.g., users, accounts, or hosts). We demonstrate that incorporating contextual information with the proposed CWAE model enables more robust anomaly detection across diverse and complex datasets, and offers significant advantages in improving detection accuracy.
@@
  • We promote a context-conditioned anomaly detection framework.
  • We propose the CWAE model to demonstrate the benefits of context-aware learning.
  • We introduce a principled method to select a suitable feature context feature for conditional modeling.
@@
  • We demonstrate that conditional models can outperform both their non-contextual counterparts of the same architecture and state-of-the-art (SOTA) unconditional models.
2. Related Works
@@
Anomaly detection has been a long-standing area of research, with early work dominated by statistical and distance-based approaches. Classical statistical methods such as Principal Component Analysis (PCA) (Shyu et al., 2003) and distance-based techniques like k-Nearest Neighbors (Liao and Vemuri, 2002) and clustering-based methods (Münz et al., 2007) aim to identify points that deviate from expected norms in the data distribution. While effective in lower dimensions, these techniques struggle in high-dimensional settings due to the “curse of dimensionality.”
@@
More recent advances leverage deep neural networks to learn expressive representations for anomaly detection in complex data. Notable lines of work include one-class classification with deep feature embeddings (Ruff et al., 2018, 2021), autoencoder-based approaches that reconstruct normal samples (Gong et al., 2019; Zhou and Paffenroth, 2017; Zhao et al., 2017; Liu et al., 2021), and generative models like GANs (Schlegl et al., 2017; Sabuhi et al., 2021), normalizing flows (Gudovskiy et al., 2022), and diffusion models (Livernoche et al., 2023). However, several studies (Kirichenko et al., 2020; Zhang et al., 2021) have shown that generative models trained to match the marginal data distribution can fail to identify out-of-distribution or context-specific anomalies, as they may assign high likelihood to atypical but structurally plausible inputs.
@@
To address this, self-supervised learning methods (Golan and El-Yaniv, 2018; Wang et al., 2019b; Chen et al., 2020; Tack et al., 2020) recast anomaly detection as an auxiliary classification task, such as predicting data augmentations or transformations. While effective in image domains, extending such techniques to tabular or categorical data is non-trivial (Bergman and Hoshen, 2020; Qiu et al., 2021; Shenkar and Wolf, 2022), and performance can degrade when the auxiliary task is misaligned with anomaly structure.
@@
Orthogonal to these, subspace anomaly detection methods aim to discover subsets of features or projections where anomalies become more apparent. Classical examples include axis-parallel subspace approaches (Kriegel et al., 2009), projection pursuit, and local subspace analysis (Aggarwal and Yu, 2013; Sathe and Aggarwal, 2016). These methods, including randomized hashing (Sathe and Aggarwal, 2016), avoid modeling the full joint distribution by seeking dimensions that isolate outliers. However, they typically treat all features equally and do not distinguish between context and behavioral attributes.
@@
In contrast, context-aware or conditional anomaly detection explicitly models how behavioral features deviate from expected values given a specific context (Song et al., 2007; Stein et al., 2020). For instance, CADENCE (Amin et al., 2019) models the conditional probability of discrete events within high-cardinality logs, and cross-linked VAEs (Shulman et al., 2020) disentangle context and content in anomaly scoring. Similarly, kernel mixture networks (Kim and Kim, 2023) estimate flexible context-conditioned distributions in sensor data. These approaches highlight the benefits of modeling conditional distributions over global joint distributions in heterogeneous environments, but they often assume that the context is known a priori, or restrict themselves to domains like time-series or event logs.
@@
Our work builds on this literature in three key ways. First, we target tabular data—a setting with less structure and no predefined context—making the task of identifying and modeling contextual dependencies significantly more challenging. Second, unlike methods that rely on dual encoders or complex density estimators, we propose a lightweight yet effective Conditional Wasserstein Autoencoder (CWAE) that models $P(\text{content}|\text{context})$ using a simple reconstruction loss and MMD regularization. Third, we introduce a practical, unsupervised context selection strategy based on early validation loss, allowing our method to automatically discover useful contextual features without supervision. We show that this strategy improves anomaly detection performance across a diverse benchmark suite, demonstrating that meaningful contextual structure often exists—even when not explicitly labeled—and can be effectively leveraged with minimal complexity.
3. Preliminaries
@@
3.1. Context-Conditional Anomaly Detection
To address the heterogeneous nature of the problem, we hypothesize that a subset of features—referred to as optimal context features—captures the greatest diversity of anomaly patterns across their unique values, thereby defining distinct contexts. For simplicity, we initially focus on a single feature that exhibits this maximal diversity, denoted as the context feature, while treating the remaining features as content features to represent anomaly patterns. This choice does not affect the subsequent theoretical results, which naturally extend to multiple context features.
@@ 3.1. Context-Conditional Anomaly Detection
This approach allows the model to more accurately profile the true boundaries of normal behavior by accounting for contextual variability, rather than relying on a single, global threshold.
4. On the Advantages of Contextual Modeling
@@
Conceptually, our context-conditioned approach parallels discriminative logic in the classic generative–discriminative framework. While generative methods strive to capture the joint distribution of inputs and labels (analogous to content and context in our framework), they can be error-prone if assumptions about the data-generating process are misspecified. By contrast, a conditional or discriminative approach sidesteps modeling the context distribution altogether, focusing solely on how content depends on context. This distinction is particularly salient when context variables (e.g., userIDs) have high cardinality, since modeling $P(\mathbf{C})$ or the full joint distribution $P(\mathbf{Y},\mathbf{C})$ becomes prohibitively complex, risking a curse of dimensionality.
@@
Such complexity motivates Vapnik’s principle (Vapnik, 1995), which asserts that one should not solve a more general, and potentially more difficult, problem than necessary to meet the ultimate predictive goal. By focusing on $P(\mathbf{Y}\mid\mathbf{C})$, our method reduces the risk of model misspecification surrounding context variables and streamlines parameter estimation in high-dimensional spaces. In practice, this can lead to more robust and data efficient performance, which is an advantage also highlighted by extensive generative vs. discriminative comparisons in the classification literature. We illustrate this in more detail in the following subsections.
@@
Furthermore, we illustrate the advantage of conditional modeling from the perspective of variance reduction by the conditioned context variable. We argue that with an optimal choice of context variable, the conditional variance given each context would be smaller, resulting in less spread among data points, making deviations or anomalies more pronounced and easier to detect.
4.1. Model Misspecification Risks
@@
Several theoretical and empirical studies substantiate these advantages for conditional or discriminative approaches. (Ng and Jordan, 2001) provides a rigorous comparison of Logistic Regression versus Naïve Bayes, demonstrating that although the generative model can dominate in very low-data regimes (due to stronger assumptions reducing variance), the discriminative model typically surpasses it once the sample size becomes sufficiently large. This phenomenon occurs precisely because a discriminative model directly optimizes the conditional likelihood pertinent to the task—akin to estimating $P(\mathbf{Y}\mid\mathbf{C})$ rather than the full joint distribution. The asymptotic analysis by (Liang and Jordan, 2008) further clarifies how misspecification in modeling $P(\mathbf{C})$ can degrade a generative model’s performance, while a discriminative estimator that only captures $P(\mathbf{Y}\mid\mathbf{C})$ remains robust. (Lasserre et al., 2006) shows how purely generative assumptions can degrade classification accuracy when they are violated, and how incorporating a discriminative component can improve robustness and prediction quality.
4.2. Curse of Dimensionality
@@
In the high-dimensional context space, (Hastie et al., 2009) discusses how focusing on the conditional distribution helps avoid unnecessary parameter explosion, while (Bishop, 2006) and (Murphy, 2012) provide extensive treatments of both paradigms, showing that discriminative techniques can be more robust to structural mismatches in the data generation process. These results generalize beyond simple supervised learning to broader scenarios such as one-class classification or semi-supervised training, where focusing on the part of the distribution necessary for the objective—be it anomaly detection or prediction—can reduce parameter complexity and improve predictive accuracy. Hence, in contexts where one can define a meaningful ”background” subset of variables, instead of learning the full joint distribution, conditioning on them directly can provide both practical and theoretical advantages in model efficiency. By extension, in high-dimensional scenarios, focusing solely on $P(\mathbf{Y}\mid\mathbf{C})$ avoids the exponential blow-up in complexity that can come with modeling or sampling from $P(\mathbf{C})$. Consequently, conditional modeling can confer both practical and theoretical gains in efficiency and accuracy, making it a compelling strategy whenever a meaningful separation between context and content variables is possible.
4.3. Variance Reduction
@@
5. Modeling Conditional Distribution with CWAE
In this work, we introduce the Contextual Wasserstein Auto-Encoder (CWAE), a straightforward autoencoder-based method for modeling context-conditional distributions in anomaly detection. CWAE is designed as a proof-of-concept to showcase how incorporating contextual information can improve detection performance. Although not intended as an optimized or state-of-the-art solution, it offers a clear and interpretable framework for illustrating the advantages of context-aware learning in anomaly detection.
5.1. CWAE Model Structure
@@ 5.1.1. Model Objective
5.1.2. Model Architecture
Figure 1. Architecture of the CWAE model utilized in the experiments.
CWAE consists of three main components: embedding layers, an encoder, and a decoder.

### v2

# Contextual Learning for Anomaly Detection in Tabular Data

@@ Abstract
Anomaly detection is critical in domains such as cybersecurity and finance, especially when working with large-scale tabular data. Yet, unsupervised anomaly detection—where no labeled anomalies are available—remains challenging because traditional deep learning methods model a single global distribution, assuming all samples follow the same behavior. In contrast, real-world data often contain heterogeneous contexts (e.g., different users, accounts, or devices), where globally rare events may be normal within specific conditions. We introduce a contextual learning framework that explicitly models how normal behavior varies across contexts by learning conditional data distributions $P(\mathbf{Y}\mid\mathbf{C})$ rather than a global joint distribution $P(\mathbf{X})$. The framework encompasses (1) a probabilistic formulation for context-conditioned learning, (2) a principled bilevel optimization strategy for automatically selecting informative context features using early validation loss, and (3) theoretical grounding through variance decomposition and discriminative learning principles. We instantiate this framework using a novel conditional Wasserstein autoencoder as a simple yet effective model for tabular anomaly detection. Extensive experiments across eight benchmark datasets demonstrate that contextual learning consistently outperforms global approaches—even when the optimal context is not intuitively obvious—establishing a new foundation for anomaly detection in heterogeneous tabular data.
1 Introduction
@@
Unsupervised anomaly detection (AD) is a critical field focused on identifying patterns that deviate from expected behavior, with applications spanning a wide range of tasks like security, fraud detection, and system reliability chandola2009anomaly; aggarwal2017outlier. The core challenge is not merely detecting new or unseen objects, but discerning whether such observations align with the learned notion of “normal,” or truly represent anomalies.
@@
Recent advances in deep learning have driven a wave of neural network-based methods for unsupervised anomaly detection landauer_deep_2023; pang2021deep. Despite methodological differences, most deep approaches model the joint distribution of all features, implicitly assuming that data patterns are homogeneous across conditions. In practice, patterns often vary significantly with context. For example, in cybersecurity, a spike in data transfer may be benign for a user who routinely performs backups but suspicious for a database administrator. Ignoring such context can degrade detection accuracy.
@@
To address this challenge, we argue probabilistically and demonstrate empirically the effectiveness of modeling conditional distributions over joint distributions for anomaly detection. We advocate for a contextual learning framework, which models conditional distributions to capture intra-context regularities while isolating inter-context variability, thereby enhancing detection performance without requiring separate models for each condition.
@@
Selecting an appropriate context feature is a critical aspect of contextual learning. It is often not obvious what feature one should condition on when modeling the data. Indeed, the choice of conditioning variable can significantly influence model performance. To address this, we develop a bi-level optimization methodology to automatically identify the context features based on early validation loss. Requiring only a small set of validation samples, our approach is simple, model-agnostic, and demonstrates consistent benefits across a wide range of datasets.
@@
To demonstrate the framework’s utility and empirically validate our claim, we build upon the Wasserstein autoencoder (WAE) tolstikhin_wasserstein_2019 to develop the conditional Wasserstein autoencoder (CWAE), a novel conditional variant designed to model conditional distributions for anomaly detection. We instantiate the framework using CWAE to conduct experiments using several tabular AD benchmark datasets containing inherent contextual information (e.g., users, accounts, or hosts). We demonstrate that incorporating contextual information with CWAE enables more robust anomaly detection across diverse and complex datasets, and offers significant advantages in improving detection accuracy.
@@
  • We introduce a general contextual learning framework for anomaly detection.
  • We provide a probabilistic formulation and theoretical justification grounded in variance decomposition and discriminative learning.
  • We propose a bilevel optimization strategy for context selection using early validation loss.
  • We instantiate the framework with a novel CWAE model and demonstrate consistent performance gains across diverse datasets.
@@
  • We demonstrate that conditional models in general can outperform both their non-contextual counterparts of the same architecture and state-of-the-art (SOTA) unconditional models.
2 Related Works
@@
Anomaly detection has been a long-standing area of research, with early work dominated by statistical and distance-based approaches. Classical statistical methods such as Principal Component Analysis (PCA) shyu2003novel and distance-based techniques like k-Nearest Neighbors liao2002use and clustering-based methods munz2007traffic aim to identify points that deviate from expected norms in the data distribution. While effective in lower dimensions, these techniques struggle in high-dimensional settings due to the “curse of dimensionality."
@@
More recent advances leverage deep neural networks to learn expressive representations for anomaly detection in complex data. Notable lines of work include one-class classification with deep feature embeddings ruff_deep_2018; ruff2021unifying, autoencoder-based approaches that reconstruct normal samples gong2019memorizing; zhou2017anomaly; zhao2017spatio; liu_rca_2021, and generative models like GANs schlegl2017unsupervised; sabuhi2021applications, normalizing flows gudovskiy2022cflow, and diffusion models livernoche2023diffusion. However, several studies kirichenko2020normalizing; zhang2021understanding have shown that generative models trained to match the marginal data distribution can fail to identify out-of-distribution or context-specific anomalies, as they may assign high likelihood to atypical but structurally plausible inputs.
@@
To address this, self-supervised learning methods golan2018deep; wang2019effective; chen2020simple; tack2020csi recast anomaly detection as an auxiliary classification task, such as predicting data augmentations or transformations. While effective in image domains, extending such techniques to tabular or categorical data is non-trivial bergman_classification-based_2020; qiu_neural_2022; shenkar_anomaly_2022, and performance can degrade when the auxiliary task is misaligned with anomaly structure.
@@
Orthogonal to these, subspace anomaly detection methods aim to discover subsets of features or projections where anomalies become more apparent. Classical examples include axis-parallel subspace approaches kriegel2009outlier, projection pursuit, and local subspace analysis aggarwal2013outlier; sathe2016subspace. These methods, including randomized hashing sathe2016subspace, avoid modeling the full joint distribution by seeking dimensions that isolate outliers. However, they typically treat all features equally and do not distinguish between context and behavioral attributes.
@@
In contrast, context-aware or conditional anomaly detection explicitly models how behavioral features deviate from expected values given a specific context song2007conditional; stein2020unsupervised. For instance, CADENCE amin2019cadence models the conditional probability of discrete events within high-cardinality logs, and cross-linked VAEs shulman2020crosslinked disentangle context and content in anomaly scoring. Similarly, kernel mixture networks kim2023conditional estimate flexible context-conditioned distributions in sensor data. These approaches highlight the benefits of modeling conditional distributions over global joint distributions in heterogeneous environments, but they often assume that the context is known a priori, or restrict themselves to domains like time series or event logs.
@@
Our work builds on this literature in three key ways. First, we target tabular data—a setting with less inherent structure and no predefined context—making the identification and modeling of contextual dependencies significantly more challenging. Second, unlike prior methods that rely on dual encoders or complex density estimators, we propose a lightweight yet effective CWAE that models $P(\text{content}\mid\text{context})$ through a simple reconstruction objective with MMD regularization. Third, we introduce a practical, unsupervised context selection strategy based on early validation loss, allowing the framework to automatically discover informative context features without supervision. Together, these components form a unified contextual learning framework for anomaly detection that is both conceptually simple and broadly applicable. This perspective reveals that meaningful contextual structure often exists even when not explicitly labeled, and that leveraging such structure can significantly enhance detection performance with minimal architectural complexity.
3 Preliminaries
@@
3.1 Contextual Learning Formulation
To address the heterogeneous nature of the problem, we hypothesize that a subset of features—referred to as optimal context features—captures the greatest diversity of anomaly patterns across their unique values, thereby defining distinct contexts. For simplicity, interpretability, and computational stability, we initially focus on a single feature that exhibits this maximal diversity, denoted as the context feature, while treating the remaining features as content features that characterize anomaly patterns within each context.
Focusing on a single feature offers several practical advantages. First, it enhances interpretability—practitioners can directly observe which variable defines the contextual partition. Second, it avoids the combinatorial explosion of context groups that arises under multi-feature conditioning, which can lead to data sparsity and unstable estimation when certain feature combinations are rarely observed. Third, our empirical findings indicate that well-chosen single features often capture the dominant axis of contextual variation, while remaining dependencies can be effectively modeled through the content features. Importantly, this simplification does not restrict the generality of our framework: all theoretical results, including those on variance decomposition and discriminative learning advantages, naturally extend to the multi-feature case where $\mathbf{C}\in\mathbb{R}^{k}$ with $k>1$. Future work may explore for multi-feature context construction when single-feature conditioning proves insufficient.
@@ 3.1 Contextual Learning Formulation
This approach allows the model to more accurately profile the true boundaries of normal behavior by accounting for contextual variability, rather than relying on a single, global threshold. This formulation constitutes the core of our contextual learning framework.
4 On the Advantages of Contextual Learning
@@
Conceptually, our contextual learning paradigm—defined as conditional modeling of content given context—parallels discriminative logic in the classic generative-discriminative framework. While generative methods strive to capture the joint distribution of inputs and labels (analogous to content and context in our framework), they can be error-prone if assumptions about the data-generating process are misspecified. By contrast, a conditional or discriminative approach sidesteps modeling the context distribution altogether, focusing solely on how content depends on context. This distinction is particularly salient when context variables (e.g., userIDs) have high cardinality, since modeling $P(\mathbf{C})$ or the full joint distribution $P(\mathbf{Y},\mathbf{C})$ becomes prohibitively complex, risking a curse of dimensionality.
@@
Such complexity motivates Vapnik’s principle (vapnik1995nature), which asserts that one should not solve a more general, and potentially more difficult, problem than necessary to meet the ultimate predictive goal. By focusing on $P(\mathbf{Y}\mid\mathbf{C})$, our method reduces the risk of model misspecification surrounding context variables and streamlines parameter estimation in high-dimensional spaces. In practice, this can lead to more robust and data efficient performance, which is an advantage also highlighted by extensive generative vs. discriminative comparisons in the classification literature. We illustrate this in more detail in the following subsections.
@@
Furthermore, we illustrate the advantage of contextual learning from the perspective of variance reduction by the conditioned context variable. We argue that with an optimal choice of context variable, the conditional variance given each context would be smaller, resulting in less spread among data points, making deviations or anomalies more pronounced and easier to detect.
4.1 Model Misspecification Risks
@@
Several theoretical and empirical studies substantiate these advantages for conditional or discriminative approaches. andrew_comparison provides a rigorous comparison of Logistic Regression versus Naïve Bayes, demonstrating that although the generative model can dominate in very low-data regimes (due to stronger assumptions reducing variance), the discriminative model typically surpasses it once the sample size becomes sufficiently large. This phenomenon occurs precisely because a discriminative model directly optimizes the conditional likelihood pertinent to the task—akin to estimating $P(\mathbf{Y}\mid\mathbf{C})$ rather than the full joint distribution. The asymptotic analysis by (liang2008asymptotic) further clarifies how misspecification in modeling $P(\mathbf{C})$ can degrade a generative model’s performance, while a discriminative estimator that only captures $P(\mathbf{Y}\mid\mathbf{C})$ remains robust. lasserre2006hybrid shows how purely generative assumptions can degrade classification accuracy when they are violated, and how incorporating a discriminative component can improve robustness and prediction quality.
4.2 Curse of Dimensionality
@@
In the high-dimensional context space, hastie2009elements discusses how focusing on the conditional distribution helps avoid unnecessary parameter explosion, while bishop2006pattern and murphy2012machine provide extensive treatments of both paradigms, showing that discriminative techniques can be more robust to structural mismatches in the data generation process. These results generalize beyond simple supervised learning to broader scenarios such as one-class classification or semi-supervised training, where focusing on the part of the distribution necessary for the objective—be it anomaly detection or prediction—can reduce parameter complexity and improve predictive accuracy. Hence, in contexts where one can define a meaningful "background" subset of variables, instead of learning the full joint distribution, conditioning on them directly can provide both practical and theoretical advantages in model efficiency. By extension, in high-dimensional scenarios, focusing solely on $P(\mathbf{Y}\mid\mathbf{C})$ avoids the exponential blow-up in complexity that can come with modeling or sampling from $P(\mathbf{C})$. Consequently, conditional modeling can confer both practical and theoretical gains in efficiency and accuracy, making it a compelling strategy whenever a meaningful separation between context and content variables is possible.
4.3 Variance Reduction
@@
5 Instantiating Contextual Learning with CWAE
To demonstrate the contextual learning framework in practice, we instantiate it using the Contextual Wasserstein Auto-Encoder (CWAE). CWAE is a novel, generative, and reconstruction-based model designed to be straightforward and interpretable, operationalizing the framework by modeling context-conditional distributions for anomaly detection. While CWAE achieves strong performance, its primary purpose is to illustrate the effectiveness and generality of the contextual learning framework rather than to serve as the main contribution of this work. Comprehensive implementation details are provided in the Appendix A.
5.1 CWAE Model Structure
@@ 5.1.1 Model Objective
5.1.2 Model Architecture
As displayed in Figure 1, CWAE consists of three main components: embedding layers, an encoder, and a decoder.
Figure 1: Architecture of the CWAE model utilized in the experiments.